# Imagine All The Relevance: Scenario-Profiled Indexing with Knowledge Expansion for Dense Retrieval

### v1

@@ 4.1 Experimental settings
Implementation details.  We construct the scenario-augmented training dataset by randomly sampling 300 documents per dataset from StackExchange split of BRIGHT (Su et al., 2024) and all datasets in BEIR (Thakur et al., 2021), resulting in a total of 8,100 documents. Note that there is not any exposure to test queries or their relevance annotations, ensuring that retrieval evaluation remains entirely independent of the training process. For each sampled document, we use GPT-4o to generate scenarios, ensuring that the dataset contains high-quality scenarios that facilitate deeper reasoning over a document’s contents. For the scenario generator, we use Llama-3.2-3B-Instruct (Dubey et al., 2024) as the backbone model and fine-tune it using LoRA (Hu et al., 2021). For our main experiments, we set the relevance weight in Equation (3) to 0.7. More details about this part are provided in Appendix C.2.
@@ 5 Analysis
Ablation study on relevance weight.  Figure 6 also illustrates the impact of the relevance weight $\alpha$ in Equation (3), which balances scenario-level and document-level relevance scores. Across nearly all components, retrieval performance peaked at $\alpha=0.7$. However, different document type exhibited distinct trends. In natural language and code documents, M+E performance improves as $\alpha$ decreases indicating that the information provided by scenarios is more crucial than the original document content in these document types. Conversely, in math documents, performance improves steadily as $\alpha$ increases, suggesting that document-level content (e.g., LaTeX equations) is more critical for effective retrieval. Given these analysis, we select $\alpha=0.7$ as the primary configuration, as it provides the most consistent performance improvements across different types of documents.
Zero-shot generalization of scenario generator.  To evaluate the generalizability of our scenario generator, we conduct a zero-shot experiment where the generator was trained exclusively on the BEIR corpus without exposure to the BRIGHT corpus. Table 2 presents the retrieved results under this setting. Notably, BEIR does not fully cover the range of domains present in BRIGHT, particularly code and math. Despite this domain gap, the version of SPIKE trained only on the BEIR corpus consistently improves retrieval performance across nearly all datasets in BRIGHT. Specifically, we observe substantial gains in code domain and TheoQ. dataset, which are not covered by the BEIR corpus. These results demonstrate the strong out-of-domain generalization capability of our scenario-based retrieval framework.
@@
+SPIKE 22.0 29.0 19.0 25.5 22.0 17.6 16.2 27.9 9.6 6.2 22.7 26.7 20.4
Table 2: Retrieval performance when scenarios are generated by a scenario generator trained without the BRIGHT corpus. We report nDCG@10 for all datasets. Avg. denotes the average score across 12 datasets and Improv. denotes the improvement rate of the average score. The best score on each model is shown in bold.
@@ 6 Conclusion
This paper proposes SPIKE, a novel dense retrieval framework designed to enhance retrieval effectiveness by explicitly modeling implicit relevance. By reframing documents into structured retrieval scenarios, SPIKE addresses the limitations of existing dense retrieval models that struggle with reasoning-intensive tasks. Our extensive experiments demonstrate that SPIKE not only enhances retrieval performance across various domains, including natural language, code, and math, but also improves usability in real-world retrieval systems and serves as an effective augmentation for RAG-based applications.

### v2

@@ 3.3 Retrieval with scenario
3.4 Efficient retrieval strategy
While SPIKE’s scenario-profiled index enhances retrieval effectiveness, a naive implementation could introduce significant latency by scoring every scenario for all documents. Such an approach would cause additional overhead and latency to scale proportionally with the corpus size ($N$). To ensure practical efficiency, we employ an efficient retrieval strategy. For a given query, SPIKE first identifies a candidate set of top-k’ documents using only the document scores ($r_{d}$). Subsequently, scenario scores ($r_{s}$) are computed exclusively for this limited subset. This approach ensures that the additional computation is bounded by the hyperparameter $k^{\prime}$ (where $k<k^{\prime}\ll N$) and does not scale with the corpus size.
@@ 4.1 Experimental settings
Implementation details.  We construct the scenario-augmented training dataset by randomly sampling 300 documents per dataset from StackExchange split of BRIGHT (Su et al., 2024) and all datasets in BEIR (Thakur et al., 2021), resulting in a total of 8,100 documents. Note that there is not any exposure to test queries or their relevance annotations, ensuring that retrieval evaluation remains entirely independent of the training process. For each sampled document, we use GPT-4o to generate scenarios, ensuring that the dataset contains high-quality scenarios that facilitate deeper reasoning over a document’s contents. For the scenario generator, we use Llama-3.2-3B-Instruct (Dubey et al., 2024) as the backbone model and fine-tune it using LoRA (Hu et al., 2021). For our main experiments, we set the relevance weight in Equation (3) to 0.7. We use the efficient retrieval strategy mentioned in Section 3.4 for all experiments, setting the value $K^{\prime}$ at 1000. For reproducibility, we also provide more details about implementation details in Appendix C.2.
@@ 5 Analysis
Ablation study on relevance weight.  Figure 6 also illustrates the impact of the relevance weight $\alpha$ in Equation (3), which balances scenario-level and document-level relevance scores. Across nearly all components, retrieval performance peaked at $\alpha=0.7$. However, different document type exhibited distinct trends. In natural language and code documents, M+E performance improves as $\alpha$ decreases indicating that the information provided by scenarios is more crucial than the original document content in these document types. Conversely, in math documents, performance improves steadily as $\alpha$ increases, suggesting that document-level content (e.g., LaTeX equations) is more critical for effective retrieval. Given these analysis, we select $\alpha=0.7$ as the primary configuration, as it provides the most consistent performance improvements across different types of documents. The full results of the ablation study on scenario components and relevance weight are provided in Appendix B.3.
Zero-shot generalization of scenario generator.  To evaluate the generalizability of our scenario generator, we conduct a zero-shot experiment where the generator was trained exclusively on the BEIR corpus without exposure to the BRIGHT corpus. Table 2 presents the retrieved results under this setting. Notably, BEIR does not fully cover the range of domains present in BRIGHT, particularly code and math. Despite this domain gap, the version of SPIKE trained only on the BEIR corpus consistently improves retrieval performance across nearly all datasets in BRIGHT. Surprisingly, even when trained solely on BEIR, it sometimes exhibits higher performance than when utilizing a scenario generator trained on the BRIGHT corpus (refer to Table 1). Specifically, we observe substantial gains in code domain and TheoT. dataset, which are not covered by the BEIR corpus. These results demonstrate the strong out-of-domain generalization capability of our scenario-based retrieval framework.
@@
+SPIKE ($C=4000$) 22.0 29.0 19.0 25.5 22.0 17.6 16.2 27.9 9.6 6.2 22.7 26.7 20.4 11.5%
+SPIKE ($C=20000$) 25.0 34.3 22.8 29.9 22.4 16.6 16.1 27.8 24.2 7.1 23.3 27.5 23.1 26.3%
[…]
+SPIKE ($C=4000$) 23.0 27.6 18.4 20.2 19.5 18.3 15.8 31.5 17.5 8.9 25.6 25.9 21.0 0.7%
+SPIKE ($C=20000$) 30.4 34.9 21.7 24.3 21.1 18.6 17.0 29.4 24.4 8.8 25.3 27.0 23.6 13.0%
[…]
+SPIKE ($C=4000$) 31.3 40.4 22.8 24.4 24.5 16.2 22.8 24.6 11.1 13.6 27.8 33.3 24.4 0.5%
+SPIKE ($C=20000$) 34.6 42.9 23.2 30.6 28.1 19.4 24.1 25.6 27.0 13.5 28.6 34.0 27.6 13.8%
Table 2: Retrieval performance when scenarios are generated by a scenario generator trained only on the BEIR corpus without the BRIGHT corpus. We report nDCG@10 for all datasets. Avg. denotes the average score across 12 datasets and Improv. denotes the improvement rate of the average score. $C$ denotes the number of documents used in the scenario-augmented training dataset. The best score on each model is shown in bold.
Ablation study on scenario-augmented training dataset.  To investigate the impact of scaling the scenario-augmented training dataset on retrieval performance, we conducted an ablation study on the size of the scenario-augmented training dataset used for training our scenario generator. Table 2 presents the results of this ablation study performed solely with the BEIR corpus. Even when utilizing only 4,000 documents, SPIKE demonstrates superior performance compared to the baseline without SPIKE. Furthermore, scaling the scenario-augmented training dataset yields a significantly greater performance improvement. Notably, as previously mentioned, this performance surpasses the retrieval performance achieved when trained on the BRIGHT dataset. These results indicate that performance of SPIKE consistently improves as the scenario-augmented training dataset scales, suggesting that SPIKE could benefit from even larger and more diverse datasets to further enhance its capabilities to capture implicit relevance across various domains.
Ablation study on efficient retrieval strategy.  To validate the effectiveness of the efficient retrieval strategy introduced in Section 3.4, we conducted an ablation studys on the candidate set size, $k^{\prime}$. Table 3 shows the retrieval performance as $k^{\prime}$ is varied. The results indicate that performance remains nearly identical to the naive method (i.e., scoring all scenarios) as long as $k^{\prime}$ is sufficiently large, while significantly reducing the computational load. This result demonstrates that our efficient retrieval strategy successfully provides high efficiency without compromising the retrieval effectiveness of the SPIKE framework.
Natural language Code Math Avg.
Bio. Earth. Econ. Psy. Sus. Rob. Stack. Leet. Pony Aops TheoQ. TheoT.
E5-Mistral 18.8 26.0 15.5 15.8 18.5 16.4 9.8 28.7 4.8 7.1 26.1 26.8 17.9
+SPIKE ($k^{\prime}=1000$) 25.9 33.0 18.2 20.6 20.6 18.4 16.2 29.4 17.5 7.0 23.4 28.4 21.6
+SPIKE ($k^{\prime}=10000$) 25.7 33.0 18.2 20.6 20.6 18.4 16.2 29.4 17.2 7.0 23.5 28.2 21.5
+SPIKE ($k^{\prime}=N$) 25.7 33.0 18.2 20.6 20.6 18.4 16.2 29.4 17.2 7.0 23.5 28.2 21.5
Table 3: Retrieval performance on BRIGHT for different values of $K^{\prime}$, using E5-Mistral-7B as the retrieval model. Retrieval performance is virtually unchanged regardless of the value of $K^{\prime}$, whether $K^{\prime}$ is set to 1000, 10000, or $N$ (i.e., exhaustive search).
@@ 6 Conclusion
This paper proposes SPIKE, a novel dense retrieval framework designed to enhance retrieval effectiveness by explicitly modeling implicit relevance. By reframing documents into structured retrieval scenarios, SPIKE addresses the limitations of existing dense retrieval models that struggle with reasoning-intensive tasks. Our extensive experiments demonstrate that SPIKE not only enhances retrieval performance across various domains, including natural language, code, and math, but also improves usability in real-world retrieval systems and serves as an effective context for RAG-based applications. Furthermore, our analysis showed that SPIKE exhibits robust out-of-domain generalization capabilities.
Acknowledgements
This work was supported by the IITP grants funded by the Korea government (MSIT) (No. RS-2020-II201361; RS-2024-00457882, AI Research Hub Project; IITP-2025-RS-2020-II201819).